What equation would best represents an ellipse for the given foci and co-vertices...
foci (0, +-3) co-vertices (+-5, 0)???

Answer by lwsshak3(11628)

What equation would best represents an ellipse for the given foci and co-vertices...
foci (0, +-3) co-vertices (+-5, 0)
..
This is an ellipse with a vertical major axis. Standard form for this elllipse:
(x-h)^2/b^2+(y-k)^2/a^2=1, a>b
..
center: (0,0)
c=3
c^2=9
b=5
b^2=25
c^2=a^2-b^2
a^2=c^2+b^2=9+25=34
a√34
Equation:
x^2/25+y^2/34=1